7: Scheduler – An organization can potentially work on multiple projects at any given time. Each

project consists of multiple tasks, which is the work that is to be performed in the project. You have to

implement a class which takes the task information from the user for a project. Then calculate completion

time of the project.

PART A: Calculate the completion time of the project given the task information. How to determine project completion time?

PART B: Consider that now you have to make the same schedule but also have to pick resources and assign

to the tasks. Each task has a special skill required for its completion (we will use 4 skill types ‘A’, ‘B’, ‘C’,

and, ‘D’). Now you will have to generate completion time by assigning resources to the tasks. If a resource

is not available, then the project will be delayed. Enter data for 10 resources using the following structure.


1
Expert's answer
2021-05-07T00:03:55-0400

1) The completion time of the project is 22 weeks based on the critical path A-C-F-G-H

2) The critical path is A-C-F-G-H

The second longest chain is A-C-E-G-H which is 21 weeks

Reduction by 3 weeks

a) Reduce A by 1 week, new project duration is 21 weeks. Crash Cost = 400

Reduce C by 2 week, Crash cost = 400 + 500*2 = 1400

Reduce by 5 weeks

a) Reduce A by 1 week, new project duration is 21 weeks. Crash Cost = 400

Reduce C by 2 week, Crash cost = 400 + 500*2 = 1400

Then at this point project duration is 19 days

Path ACFGH= 19 weeks

ACEGH = 18 weeks; ABEGH= 18 weeks; ADFGH = 18 weeks

Crash G. Crash Cost = 400 + 500*2 + 300 = 1700

Path ACFGH= 18 weeks

ACEGH = 17 weeks; ABEGH= 17 weeks; ADFGH = 17 weeks

Crash F Crash Cost =  400 + 500*2 + 300 + 1000 = 2700

Path ACFGH= 17 weeks

ACEGH = 17 weeks; ABEGH= 17 weeks; ADFGH = 16 weeks

The minimum time duration is 17 weeks